Area of a rug. Find the trinomial A(x) that represents the area of a rectangle rug whose sides are x + 3 and 2x – 1 feet. Find A(4).
Answer by mananth(16946)

A(x)=(x+3)(2x-1)
= 2x^2+5x-3
A(4) = 2*4^2+5*4-3
=32+20-3
=49
...
m.ananth@hotmail.ca